# On The Topology of Polygonal Meshes

### v1

@@ Abstract
This paper is an introductory and informal exposition on the topology of polygonal meshes. We begin with a broad overview of topological notions and discuss how homeomorphisms, homotopy, and homology can be used to characterize topology. We move on to define polygonal meshes and make a distinction between intrinsic topology and extrinsic topology which depends on the embedding space. A distinction is also made between quantitative topological properties and qualitative properties. We outline a proof of the Euler and the Euler-Poincaré formulas. The Betti numbers are defined in terms of the Euler-Poincaré formula and other mesh statistics rather than as cardinalities of the homology groups which allows us to avoid abstract algebra. Finally, we discuss how it is possible to cut a polygonal mesh such that it becomes a topological disc.
@@ 1 Introduction
From a practical point of view, topological properties can be of great interest. For instance, we may want to count the number of objects in a 2D or 3D image; this is such a common task that it is barely recognized as being related to topology. In other cases we may want to determine whether objects contain cycles or have a purely branching structure, or we may be interested in knowing whether an object has internal cavities.
@@
While there are many applications of topology, it can be a challenging topic unless one is in possession of a good mathematical background; in particular, a firm grasp of abstract algebra is needed for algebraic topology. In this document, we target applications in visual computing and sidestep some of the mathematical requirements by considering only the topology of 3D surfaces provided as polygonal meshes. Moreover, we derive all quantitative topological properties from the Euler-Poincaré formula. By limiting the scope in this way, we are able to arrive at formulas for key topological properties such as genus and Betti numbers without group theory [alexandroff2012introduction]. Since Betti numbers are almost always defined in algebraic terms this document takes an unusual route which I hope is illuminating for readers who wish to get a grasp on topology without necessarily understanding homology.
@@
1.1 What is a Hole?
Topological properties often feel familiar and appeal to our intuition, but subtleties arise when we try to make the notions precise. Perhaps the most pernicious source of confusion is the fact that solid objects and surfaces are very different from a topological point of view. This is true even in the cases where the surface is exactly identical to the boundary of some solid. For instance, consider an egg-like object which is hollow inside. We might consider this object to consist of two surfaces: the outer surface and the surface that encloses the interior void. However, if we restrict our topological analysis to surfaces only, there is no way to distinguish an object with an interior cavity from two objects where one just happens to be inside the other. Since we will only be dealing with polygonal meshes in the following, it is worthwhile to be quite clear that polygonal meshes are surfaces, and there is no way of telling (from the representation itself) whether two disjoint polygonal meshes are part of the boundary of the same solid.
Another sticking point is related to the very notion of a hole. Consider a piece of paper. It is obvious that we can cut a hole in the paper with a pair of scissors. Now, consider a donut. Clearly, the donut also has a hole, but it is of a different type. Holes of the donut variety are shown in Figure 1.
@@ 1.2 Background
Two topological spaces are said to be homeomorphic if there is a continuous bijective map from one to the other and this map has a continuous inverse (cf. Figure 2 top left). This is the strictest test of topological equivalence, but given two objects of identical, topology there is no general method for computing a homeomorphism between them. Moreover, homeomorphisms only allow us to make statements about topological equivalence, they do not allow for quantification of topological properties. Nevertheless, we will often need to make statements about topological equivalence in terms of homeomorphisms in what follows. For instance, if we say that something is a topological disc, it means that there is a homeomorphism between the object and a disc.
@@
A topological space, $X$, is a 2-manifold if any sufficiently small neighborhood of a point, $\mathbf{p}\in X$, is homeomorphic to an open disc (see the margin figure on the right). Intuitively, this means that we can cut out a (possibly tiny) patch which contains $\mathbf{p}$ and which can be flattened and stretched to a perfect disc. Thus, a 2-manifold models $\mathbb{R}^{2}$ locally at every point and can be seen as a mathematical definition of the notion of a closed surface. Of course, some surfaces, e.g. a sheet of paper, are open, and an open surface is a 2-manifold with boundary. Points on the boundary are homeomorphic to half-open disc as shown on the right.
The notion of a 2-manifold generalizes beyond surfaces in 3D. An $n$-manifold in $\mathbb{R}^{d}$ where $n\leq d$ is a set of points that is locally homeomorphic to either an open $n$-ball or a half-open $n$-ball in a sufficiently small region around every point belonging to the $n$-manifold. Because we will only consider surfaces, manifolds will always refer to 2-manifolds in the following.
@@
Another way to characterize topology is through homotopy. Two topological spaces are considered homotopic if one can be continuously deformed into the other (cf. Figure 2 middle bottom). At first glance, it may appear to be the same thing for two objects to be homeomorphic and homotopic, but that is actually false. A solid sphere is homotopic to a single point since we can smoothly contract the sphere to a point, but there is clearly no way to establish a bijective continuous mapping with a continuous inverse from a solid sphere to a single point. Conversely, we can cut a torus, tie it into a knot, and then stitch the ends back up. The knotted torus will still be homeomorphic to the original, but it is not homotopic since the continuous deformation would get stuck on the knot.
@@
Given an appropriate equivalence relation, groups can be partitioned into equivalence classes called quotient groups. If we consider, $H_{1}$, the quotient group (formed by the division-like notation in Eq. (1)) should be understood as a partitioning of all closed chains of edges ($Z_{1}$) by the equivalence relation that they differ only by the boundary of a region ($B_{1}$). In the case of an object that has sphere topology, all closed chains of edges are boundaries of some region. This means that they all differ from the empty cycle by the boundary of a region and we say that such chains are homologous to zero. On a torus there are two distinct families of chains of edges that do not bound any regions. Within these families, each chain differs from every other chain belonging to the same family by the boundary of a region, but unlike on the sphere we cannot make them go away by subtracting the boundary of a region. Summing up, on topological spheres all closed chains of edges are homologous to zero, but on topological tori there are two families of chains of edges which are not homologous to zero, and this is why $|H_{1}|=2$ for the torus. See Figure 2 (top right) where the red and green curves are representatives from each of these two families.
@@ 2 Polygonal Meshes
A polygonal mesh is a tuple, $\mathcal{M}$ = $\langle\mathcal{V},\mathcal{E},\mathcal{F}\rangle$, where $\mathcal{V}$ is a set of vertices, $\mathcal{E}$ is a set of edges, and $\mathcal{F}$ is a set of faces. We will often need to refer to the number of vertices, $V=|\mathcal{V}|$, the number of edges, $E=|\mathcal{E}]$, and the number of faces, $F=|\mathcal{F}|$. Formally, we will define an edge, $e=\{u,v\}$ where $u,v\in\mathcal{V}$ and $e\in\mathcal{E}$, as a set of two vertices, and a face, $f=\{e_{0},...,e_{n}\}$ where $e_{i}\in\mathcal{E}$ and $f\in\mathcal{F}$, as a set of edges. By this definition, vertices that belong to an edge must also be in the set of vertices and likewise edges in a face must belong to the set of edges. On the other hand, we may allow “dangling” edges not associated with any face or loose vertices that do not belong to an edge, but in practice such unused entities are often considered defects.
For the mesh to be geometrically meaningful, there must be a point, $\mathbf{p}_{v}\in\mathbb{R}^{3}$, associated with each vertex, $v\in\mathcal{V}$. If this association has been made, we can think of vertices as points in 3D, of edges as line segments, and of faces as polygons.
Validity of Faces
It must be true that if a vertex, $v$, belongs to any edge, $e$, of a face, $f$, then it also belongs to precisely one other edge since that ensures that the edges form a simple, closed cycle. In other words, the set of edges that contain a given vertex while belonging to a given face is either empty or of cardinality two:
$\forall v\in\mathcal{V},\forall f\in\mathcal{F}:\left|\left\{e|v\in e\wedge e\in f\right\}\right|\in\{0,2\}\,.$ (2)
We must avoid faces whose boundary consists of more than a single cycle of edges. This could be a practical issue if we, say, want to make a tunnel through a cube. The issue is that topologically we cannot always distinguish this configuration from two separate meshes (e.g. cube and tunnel). Thus, it is nearly the same issue as when we discussed interior voids, but in this case it is fixable since we can add edges to the mesh connecting the tunnel to the cube.
Validity of Edges
We require that edges belong to either one or two faces, i.e.
$\forall e\in\mathcal{E}:\left|\{f|e\in f\wedge f\in\mathcal{F}\}\right|\in\{1,2\}\,.$ (3)
This requirement is part of ensuring that the mesh is manifold. The edges that belong to only one face are called boundary edges. Just like the edges that bound a face, boundary edges also form cycles which bound holes in the polygonal mesh. The fact that meshes may have boundary cycles means that polygonal meshes do not always represent solids. However, even polygonal meshes which have no boundary cycles (i.e. all edges are adjacent to exactly two faces) are usually considered to be empty shells rather than solids for topological purposes.
If manifoldness (cf. Section 2.3.1) is not a concern, this requirement can be omitted.
Validity of Vertices
Another manifoldness condition (cf. Section 2.3.1) pertains to vertices. The faces incident on a vertex must be connected via shared edges. Again, this can be omitted if manifoldness is not a concern.
Multiple Incidence
@@
A special case that bears considering is that of multiple incidence. An edge might be incident on the same face twice as shown in the figure on the right. Likewise, a vertex can be incident on the same face multiple times. Although multiple incidence is not problematic in itself, it can lead to problems if we manipulate a mesh in a computer. For instance, we might merge the two faces that are incident on a given edge. If the two faces are really the same, we create two disconnected cycles that bound the same face. As discussed above, this should be avoided.
@@ 2.1 Computer Representation
In the present context, the representation generally does not matter, but there is one further concern that should be addressed. While at least one popular data format stores the vertices directly as 3D points in space this is not a good idea if we are interested in topology. Vertices should be identified by indices, allowing us to easily find out if the corners of two polygons correspond to the same vertex (simply by checking whether the vertex indices are the same). Of course, if we are interested in the geometry of the mesh, we would then need a separate table that maps vertex indices to points in space.
One type of non-manifoldness is generally accepted in practice: sometimes, a vertex is incident on more than two boundary edges. An example is shown in the figure on the right where two triangles share a vertex. margin: [Uncaptioned image] The result is that the vertex is incident on four boundary edges. As long as the mesh representation keeps track of which boundary edges that are adjacent in the incident boundary cycles, it is not a problem to have any even number of boundary edges incident on a vertex, but it is still required that the faces incident on the vertex all come in connected components that contribute two boundary edges. Put differently, the configuration is considered valid if it can be made manifold by closing the holes.

### v2

@@ Abstract
This paper is an introductory and informal exposition on the topology of polygonal meshes. We begin with a broad overview of topological notions and discuss how homeomorphisms, homotopy, and homology can be used to characterise topology. We move on to define polygonal meshes and make a distinction between intrinsic topology and extrinsic topology which depends on the embedding space. A distinction is also made between quantitative topological properties and qualitative properties. Next, we outline proofs of the Euler and the Euler-Poincaré formulas. The Betti numbers are then defined in terms of the Euler-Poincaré formula and other mesh statistics rather than as cardinalities of the homology groups which allows us to avoid abstract algebra. Finally, we discuss how it is possible to cut a polygonal mesh such that it becomes a topological disc.
@@ 1 Introduction
From a practical point of view, topological properties can be of great interest. For instance, we may want to count the number of objects in a 2D or 3D image; this is such a common task that it is rarely recognised as being related to topology. In other cases we may want to determine whether objects contain cycles or have a purely branching structure, or we may be interested in knowing whether an object has internal cavities.
@@
While there are many applications of topology, it can be a challenging topic unless one is in possession of a good mathematical background; in particular, a firm grasp of abstract algebra is needed for algebraic topology. In this document, we target applications in visual computing and sidestep some of the mathematical requirements by considering only the topology of 3D surfaces provided as polygonal meshes. Moreover, we derive all quantitative topological properties from the Euler-Poincaré formula. By limiting the scope in this way, we are able to arrive at formulas for key topological properties such as genus and Betti numbers without group theory [alexandroff2012introduction]. Since Betti numbers are almost always defined in algebraic terms, this document takes an unusual route which I hope is illuminating for readers who wish to get a grasp on topology without necessarily understanding homology.
@@
1.1 Why is topology hard?
Topological properties often feel familiar and appeal to our intuition, but complications arise when we make the notions precise. One common source of confusion is related to the very notion of a hole. Consider a plastic bag that has been sealed. We could cut a hole in the bag with a pair of scissors in order to reach the inside of the bag. Now, consider a donut. Clearly, the donut already has a hole, but it is a different type of hole from the one in the plastic bag. Holes of the donut variety are shown in Figure 1.
Another pernicious source of confusion is the fact that the same object can be understood in different ways, and these different ways may lead to different conclusions regarding its topology. This will be quite relevant in the following when we disentangle intrinsic topology which is indifferent to how the object is embedded in space and extrinsic topology which does take the spatial arrangement into account.
To give another example of the importance of the framing, consider an egg-shaped object that is hollow inside. We might consider this object to consist of two surfaces: the outer surface and the surface that encloses the interior void. However, if we restrict our topological analysis to surfaces only, there is no way to distinguish an object with an interior cavity from two objects where one just happens to be inside the other. Since we will only be dealing with polygonal meshes in the following, it is worthwhile to be quite clear that polygonal meshes are surfaces, and there is no way of telling (from the representation itself) whether two disjoint polygonal meshes are part of the boundary of the same solid.
@@ 1.2 Background
Before we turn to polygonal meshes, we will discuss some fundamental ideas that are key to understanding the topology of surfaces in general. First off, we need the notion of a topological space. We will only consider topological spaces, $X\subset\mathbb{R}^{d}$, which are subsets of the normal Euclidean space and thus inherit the basic connectivity of $\mathbb{R}^{d}$. In the following, we will often refer to the topology of “shapes” or “objects”. These two words are simply used as more familiar names for topological spaces.
@@
Two topological spaces are said to be homeomorphic if there is a continuous bijective map from one to the other and this map has a continuous inverse (cf. Figure 2 top left). This is the strictest test of topological equivalence, but given two objects of identical topology there is no general method for computing a homeomorphism between them. Moreover, homeomorphisms only allow us to make statements about topological equivalence, they do not allow for quantification of topological properties. Nevertheless, we will often need to make statements about topological equivalence in terms of homeomorphisms in what follows. For instance, if we say that something is a topological disc, it means that there is a homeomorphism between the object and a disc.
@@
A topological space, $X$, is a 2-manifold if any sufficiently small neighbourhood of a point, $\mathbf{p}\in X$, is homeomorphic to an open disc (see the margin figure on the right). Intuitively, this means that we can cut out a (possibly tiny) patch which contains $\mathbf{p}$ and which can be flattened and stretched to a perfect disc. Thus, a 2-manifold models $\mathbb{R}^{2}$ locally at every point and can be seen as a mathematical definition of the notion of a closed surface. Of course, some surfaces, e.g. a sheet of paper, are open, and an open surface is a 2-manifold with boundary. Points on the boundary are homeomorphic to half-open disc as shown on the right.
The notion of a 2-manifold generalises beyond surfaces in 3D. An $n$-manifold in $\mathbb{R}^{d}$ where $n\leq d$ is a set of points that is locally homeomorphic to either an open $n$-ball or a half-open $n$-ball in a sufficiently small region around every point belonging to the $n$-manifold. Because we will only consider surfaces, manifolds will always refer to 2-manifolds in the following.
@@
Two maps, $f,g:X\rightarrow Y$, between topological spaces are considered homotopic if one can be continuously transformed into the other, and the continuous transformation, $H:X\times[0,1]\rightarrow Y$ where $H(\mathbf{x},0)=f(\mathbf{x})$ and $H(\mathbf{x},1)=g(\mathbf{x})$, is called an homotopy.
For our purposes it is enough to note that two topological spaces are considered homotopy equivalent if one can be continuously deformed into the other (cf. Figure 2 middle bottom). At first glance, it may appear to be the same thing for two objects to be homeomorphic and homotopy equivalent, but that is actually false. A solid sphere is homotopy equivalent to a single point since we can smoothly contract the sphere to a point, but there is clearly no way to establish a bijective continuous mapping with a continuous inverse from a solid sphere to a single point. Conversely, we can cut a torus, tie it into a knot, and then stitch the ends back up. The knotted torus will still be homeomorphic to the original, but it is not homotopy equivalent since the continuous deformation would get stuck on the knot.
Shapes that belong to the same equivalence class are said to be of the same homotopy type. For instance, this is the case for the sphere and the point or the annulus and the ring in Figure 2.
@@
Given an appropriate equivalence relation, groups can be partitioned into equivalence classes called quotient groups. If we consider, $H_{1}$, the quotient group (formed by the division-like notation in Eq. (1)) should be understood as a partitioning of all closed chains of edges (i.e. the members of $Z_{1}$) by the equivalence relation that they differ only by the boundary of a region (any element of $B_{1}$). In the case of an object that has sphere topology, all closed chains of edges are boundaries of some region. This means that they all differ from the empty cycle by the boundary of a region and we say that such chains are homologous to zero. On a torus there are two distinct classes of chains of edges that do not bound any regions. One of these classes contains chains that go around the tube and the other contains chains that go around the hole of the torus. See Figure 2 (top right) where the two red curves go around the tube and the green curve goes around the inner hole. The number of generators is called the rank, $|H_{1}|$, of the group. Thus, $|H_{1}|=2$ for the torus.
Given two different chains that belong to the same class there is some patch on the surface that has their combination as its boundary, and we say that the two chains differ by this boundary. For instance, in Figure 2 (top right) the two red curves bound a ribbon around the torus and differ by the boundary of this ribbon.
@@ 2 Polygonal Meshes
Figure 3: A polygonal mesh, $\mathcal{M}=\langle\mathcal{V},\mathcal{E},\mathcal{F}\rangle$, where $\mathcal{V}=\{v_{0},v_{1},v_{2},v_{3},v_{4},v_{5}\}$, $\mathcal{E}=\{e_{0},e_{1},e_{2},e_{3},e_{4},e_{5}\}$, and $\mathcal{F}=\{f_{0},f_{1}\}$. $V=5$, $E=6$. and $F=2$. The boundaries of the two faces are $\partial f_{0}=\{e_{0},e_{1},e_{2},e_{3}\}$ and $\partial f_{1}=\{e_{1},e_{4},e_{5}\}$. The edge boundaries are $\partial e_{0}=\{v_{0},v_{1}\}$, $\partial e_{1}=\{v_{1},v_{2}\}$, $\partial e_{2}=\{v_{2},v_{3}\}$, $\partial e_{3}=\{v_{0},v_{3}\}$, $\partial e_{4}=\{v_{1},v_{4}\}$, and $\partial e_{5}=\{v_{4},v_{2}\}$.
[…]
A polygonal mesh can be understood as a generalisation of the (ancient) notion of a polyhedron. Like a polyhedron, a polygonal mesh is made of polygons that are joined along the edges. However, unlike a polyhedron, a mesh need not be closed, and the polygons are not necessarily perfectly flat. Formally, a polygonal mesh is a tuple, $\mathcal{M}$ = $\langle\mathcal{V},\mathcal{E},\mathcal{F}\rangle$, where $\mathcal{V}$ is a set of vertices, $\mathcal{E}$ is a set of edges, and $\mathcal{F}$ is a set of polygonal faces. Vertices, edges, and faces are all considered mesh elements. We will often need to refer to the number of vertices, $V=|\mathcal{V}|$, the number of edges, $E=|\mathcal{E}|$, and the number of faces, $F=|\mathcal{F}|$. We can think of vertices as 0-dimensional points in 3D and of edges as 1-dimensional line segments bounded by two vertices. Finally, faces are 2-dimensional polygons bounded by edges. A simple polygonal mesh consisting of just two faces is shown in Figure 3.
2.1 Boundaries
Faces are 2-dimensional patches bounded by edges. Likewise, edges are 1-dimensional curves (usually considered line segments) bounded by two vertices. The boundary relationship is formalised using the operator, $\partial$, which maps a set of $n$-dimensional mesh elements to their boundary which is a set of (n1)(n-1)-dimensional mesh elements. The boundary of a face is a set of edges,
$\partial f=\{e_{0},...,e_{n}\}\,$, (2)
where $e_{i}\in\mathcal{E}$ and $f\in\mathcal{F}$. Likewise, the boundary of an edge is the set containing its two vertices,
$\partial e=\{u,v\}\,$, (3)
where $u,v\in\mathcal{V}$ and $e\in\mathcal{E}$. Vertices do not have a boundary, hence
$\partial v=\emptyset\,$, (4)
where $v\in\mathcal{V}$.
It is possible to talk about the boundary of a set of mesh elements, but first we need to define the (set-valued) co-boundary relation,
$\delta_{\mathcal{X}}(\epsilon)=\{x_{i}\in{\mathcal{X}}|\epsilon\in\partial x_{i}\}\,.$ (5)
[…]
For a set, $\mathcal{X}$}, of mesh elements, $\delta_{\mathcal{X}}(\epsilon)$ is the subset of elements in $\mathcal{X}$} whose boundaries contain $\epsilon$. Now, we define the boundary of $\mathcal{X}$}:
$\partial{\mathcal{X}}=\{\epsilon\in\bigcup_{x_{i}\in{\mathcal{X}}}\partial x_{i}|\left|\delta_{\mathcal{X}}(\epsilon)\right|=1\}$ (6)
In the case of faces, the boundary is thus the edges that belong to the boundary of precisely one face from the set. The analogous statement is true for a set of edges while for a set of vertices the boundary is empty.
Continuing the example from Figure 3, if we let $\mathcal{F}=\{f_{0},f_{1}\}$ denote both faces (i.e. the entire mesh), it is clear that $\partial\mathcal{F}=\{e_{0},e_{2},e_{3},e_{4},e_{5}\}$, since the shared edge $e_{1}$ is excluded from the boundary of $\mathcal{F}$.
2.2 Cycles
Figure 4: Left: two sets of edges where the left one is not a cycle, but the right one is a cycle. The cycle instigator is shown as a red line segment. Right: a cube which is not a cycle since it is missing a face is shown on the left. On the right a closed cube which constitutes a cycle. Both of the shown cycles are also simple.
A cycle is simply a set, $\mathcal{C}$, of mesh elements of the same dimension which has an empty boundary. In other words, if $\mathcal{C}$ is a cycle it is true that
$\partial\mathcal{C}=\emptyset\,.$ (7)
Sets of faces or edges are closed if they are cycles.
The notion of cycles is important in topology, and we can define cycles for sets of vertices, edges, and faces, although edge cycles are the most interesting. In fact, vertices and sets of vertices are trivially cycles.
A cycle, $\mathcal{C}$, is simple if it consists of a single connected component and fulfils the following predicate:
$\text{SimpleCycle}(\mathcal{C})\iff\forall\epsilon\in\bigcup_{x_{i}\in\mathcal{C}}\partial x_{i}:\left|\delta_{\mathcal{C}}(\epsilon)\right|=2\,.$ (8)
This means that a cycle of edges is simple if it is a connected set of edges and all vertices that are incident on the cycle are incident on precisely two edges. Note also that vertices and sets of vertices are not simple cycles.
The boundaries of faces of a polygonal mesh must be simple cycles:
$\forall f\in\mathcal{F}:\text{SimpleCycle}(\partial f)\,.$ (9)
This is an important requirement and it is a matter of definition although, in a few cases, we may want to deviate slightly from the requirement that face boundaries must be simple (cf. Section 2.4).
2.3 Manifoldness
  1. 1.
[…]
    $\forall e\in\mathcal{E}:\left|\delta_{\mathcal{F}}(e)\right|\in\{1,2\}$ (10)
  2. 2.
    The mesh elements incident on a vertex, $v$, form only a single connected component, i.e. it must be true that
    $\forall e_{0},e_{1}\in\delta_{\mathcal{E}}(v):e_{0}\sim e_{1}$ (11)
    where the transitive relation $\sim$ is defined by the conditions that
    $e_{0}\sim e_{1}$ $\impliedby$ $\delta_{\mathcal{F}}(e_{0})\cap\delta_{\mathcal{F}}(e_{1})\neq\emptyset$ (12)
    $e_{0}\sim e_{1}$ $\impliedby$ $\exists e_{2}\in\delta_{\mathcal{E}}(v):e_{0}\sim e_{2}\wedge e_{2}\sim e_{1}$ (13)
If $\mathcal{M}$ passes both of these tests, we know that it can everywhere be flattened to a disc (or half disc) and is thus manifold (possibly with boundary).
Provided a mesh is manifold we can prove that boundary edges always form cycles. As stated above, the faces incident on a vertex must form a single component. This means that a vertex is always incident on either two or zero boundary edges since the incident faces must form either a fan (as shown right) or a disc around the vertex. margin: [Uncaptioned image]
[…]
Consequently, all vertices which belong to boundaries are incident on two boundary edges (see the figure on the right). This means that we can always jump from a boundary edge to the next edge across one of its vertices, but, assuming the mesh is finite, we must eventually revisit the edge we started from.
A mesh which is produced by joining two pyramids (one upside down) such that the tips of the pyramids are a shared vertex is not manifold. We could try to flatten a region around the shared vertex, but it would always end up as two discs connected at a single point. Likewise, if two cubes are placed side by side such that two edges touch (forming a shared edge) that is also a non-manifold configuration since four faces then meet at the edge where the cubes are joined. The two non-manifold configurations are shown in Figure 5.
@@ 2.4 Computer Representation
In the present context, the representation generally does not matter, but there is one further concern that should be addressed. While at least one popular data format stores the vertices directly as 3D points in space this is not a good idea if we are interested in topology. Vertices should be identified by indices, allowing us to easily find out if the corners of two polygons correspond to the same vertex (simply by checking whether the vertex indices are the same).
Of course, if we are interested in the geometry of the mesh, we need a separate table that maps vertex indices to points in space.
Acceptable Non-manifold Configurations
One type of non-manifoldness is generally accepted in practice. Sometimes, a vertex is incident on more than two boundary edges. margin: [Uncaptioned image] An example is shown in the figure on the right where two triangles share a vertex. The result is that the vertex is incident on four boundary edges. As long as the mesh representation keeps track of which boundary edges that are adjacent in the incident boundary cycles, it is not a problem to have any even number of boundary edges incident on a vertex, but it is still required that the faces incident on the vertex all come in connected components that contribute two boundary edges. Put differently, the configuration is considered valid if it can be made manifold by closing the holes.
Multiple Incidence
[…]
A special case that bears considering is that of multiple incidence. An edge might be incident on the same face twice. Likewise, a vertex can be incident on the same face multiple times. It is probably clear that this breaks the simple-cycle requirement, but in practice mesh libraries can often handle such configurations.
Although multiple incidence is not necessarily problematic in itself, it can lead to problems if we manipulate a mesh in a computer. For instance, we might merge the two faces that are incident on a given edge. If the two faces are really the same as in the example shown on the right, we create two disconnected cycles that bound the same face. As discussed above, this should be avoided.
Triangle Meshes